### v1

# Improving Statistical Significance in Human Evaluation of Automatic Metrics via Soft Pairwise Accuracy

Abstract
Selecting an automatic metric that best emulates human judgments is often non-trivial, because there is no clear definition of “best emulates.” A meta-metric is required to compare the human judgments to the automatic metric judgments, and metric rankings depend on the choice of meta-metric. We propose Soft Pairwise Accuracy (SPA), a new meta-metric that builds on Pairwise Accuracy (PA) but incorporates the statistical significance of both the human judgments and the metric judgments. SPA allows for more fine-grained comparisons between systems than a simplistic binary win/loss, and addresses a number of shortcomings with PA: it is more stable with respect to both the number of systems and segments used for evaluation, it mitigates the issue of metric ties due to quantization, and it produces more statistically significant results. SPA was selected as the official system-level metric for the 2024 WMT metric shared task.
Improving Statistical Significance in Human Evaluation
of Automatic Metrics via Soft Pairwise Accuracy
Brian Thompsonthanks: Correspondence: brianjt@amazon.com. Work is unrelated to and conducted independently from the author’s position at Amazon. Amazon                        Nitika Mathur Oracle                        Daniel Deutsch Google                        Huda Khayrallah Microsoft
1 Introduction
Automatic metrics are crucial because researchers and practitioners in NLP typically can’t always afford the high cost and latency of high-quality human evaluations. Despite their shortcomings, metrics like Word Error Rate (WER) and BLEU Papineni et al. (2002)—in conjunction with carefully curated test sets—have been crucial for the field of NLP, as they have provided a yardstick to make continual progress over many decades in Automatic Speech Recognition (ASR) and Machine Translation (MT), respectively.
Reliance on automatic metrics makes selecting a good automatic metric of paramount importance. Conceptually, an automatic metric should emulate human judgments. Selecting an automatic metric typically entails generating a set of human judgments for a wide variety of outputs from a large number of different systems, and selecting the automatic metric that produces scores most similar to the human judgments. But what does “most similar” even mean? To select the “most similar” metric we need a meta-metric to compare metric scores to human judgments. The choice of meta-metric has been hotly debated in the field of MT (see § 5).
Somewhat recently, Kocmi et al. (2021) made a case for choosing metrics with the best pairwise rankings—i.e. Pairwise Accuracy (PA) as a meta-metric—as the typical use case for a metric is to directly compare two systems on a given test set. This has been adopted by the WMT Metrics shared task for the past several years (Freitag et al., 2021, 2022, 2023). However, standard best practice when using an automatic metric is to consider not only which system a metric prefers, but also whether or not that preference is statistically significant. Thus we argue that metrics should not only emulate the accuracy of human pairwise ranking, they should also emulate the confidence or statistical significance of the human judgments.
To this end, we propose Soft Pairwise Accuracy (SPA), which considers statistical significance of both the metric in question and the human judgments when evaluating the extent to which the metric agrees with the human judgments. We show that SPA, as its name implies, can be viewed as a soft (i.e. non-quantized) version of PA, and present analysis that demonstrates SPA has several distinct advantages over PA. In particular, we find it is more stable with respect to the number of segments and systems being scored, it does not produce a large number of ties between metrics, and it produces substantially more statistically significant comparisons between metrics. As a result of our recommendations, SPA is the official system-level metric for the 2024 WMT metric shared task.
2 Method
We propose a simple meta-metric for evaluating automatic metrics given human judgments, which we denote Soft Pairwise Accuracy:
$SPA={{N\choose 2}}^{-1}\,\,\sum_{i=0}^{N-1}\sum_{j=i+1}^{N-1}1-\lvert p_{ij}^{h}-p_{ij}^{m}\rvert$ (1)
where $N$ is the number of systems, $p_{ij}^{h}$ is the $p$-value for hypothesis that system $i$ is better than system $j$ given the human judgments, and $p_{ij}^{m}$ is the $p$-value for hypothesis that MT system $i$ is better than system $j$ given the metric scores, and the first term is normalizing by the total number of pairs: (${{N\choose 2}}^{-1}=\frac{2}{N(N{-}1)}$. The $p$-value formulation will be defined in § 3.2, but for now, intuitively it can be thought of as an approximation to the confidence of the human raters and automatic metric. In particular, $p_{ij}\approx 0$ means the raters or metric are extremely confident that system $i$ is better than system $j$, $p_{ij}\approx 1$ means that the raters or metric are extremely confident that system $j$ is better than system $i$, and $p_{ij}\approx 0.5$ means that the raters or metric believe the two systems to be indistinguishable in terms of performance.
Figure 1: Illustration of the individual components used to calculate both pairwise confidence error and pairwise accuracy for the Prism metric (Thompson and Post, 2020a, b) on the WMT 2023 English-German language pair. Each box represents a comparison between two systems, systems $i$ and $j$. MT systems are sorted by average human judgment score for easier interpretation. The right column is one minus the absolute difference between the human preference for systems $i$ over system $j$ (left column) and the metric preference for system $i$ over system $j$ (middle column). In PA (top row), human and metric preferences are quantized to 0 and 1, and PA is thus an average of binary terms. In SPA (bottom row), human and metric preferences range from 0 to 1, and as a result SPA is an average of values ranging from 0 to 1. SPA can be viewed as a "soft" extension to pairwise accuracy that incorporates both human judgment and metric uncertainty, allowing for partial credit.
2.1 Relationship to Pairwise Accuracy
Pairwise accuracy is defined as
$PA={{N\choose 2}}^{-1}\,\,\sum_{i=0}^{N-1}\sum_{j=i+1}^{N-1}\,a_{ij}^{m}$ (2)
where $\,a_{ij}^{m}$ is 1 when the metric and human judgments agree and 0 otherwise.
A $p$-value $p_{ij}$ will be less than 0.5 when the human raters (or automatic metric) prefer system $i$ over system $j$, and greater than 0.5 when the human raters (or automatic metric) prefer system $j$ over system $i$.
This allows us to define PA in terms of quantized $p$-values:
$PA={{N\choose 2}}^{-1}\,\,\sum_{i=0}^{N-1}\sum_{j=i+1}^{N-1}1-\left|\lfloor\,p_{ij}^{h}\rceil{-}\lfloor\,p_{ij}^{m}\rceil\right|$ (3)
Where rounding is denoted as:
$\lfloor x\rceil=\begin{dcases*}1&$x\geq 0.5$\\ 0&$x<0.5$\\ \end{dcases*}$
Comparing Equation 1 and Equation 3, highlights the fact that SPA can be viewed as a ‘soft’ extension to pairwise accuracy that incorporates uncertainty in both the human and metric scores. A visualization of this is provided in Figure 1. We hypothesize that this rounding/quantization is effectively acting as additive random noise on top of the underlying SPA meta-metric.
In cases where both the MT metric and the human evaluation both have high statistical significance (regardless of whether the metric agrees with the human judgments or not, i.e. $p_{ij}^{m}\approx 0$ or $p_{ij}^{m}\approx 1$), the contribution of that system pair to SPA and PA is approximately identical. However, there are two important cases where our meta-metric differs from pairwise accuracy:
  1. 1.
    The human evaluation has high significance (e.g. $p_{ij}^{h}\approx 0$ or $p_{ij}^{h}\approx 1$), but the metric has low statistical significance (e.g. $p_{ij}^{m}\approx 0.5$): Even if the metric happens to choose the correct winner, we partially penalize the metric for not having high statistical significance.
  2. 2.
    The human evaluation finds the systems are approximately tied (e.g. $p_{ij}^{h}\approx 0.5$): In this case, we partially penalize the metric if has high statistical significance (e.g. $p_{ij}^{m}\approx 0$ or $p_{ij}^{m}\approx 1$) even if it happens to pick the same winner as the human evaluation, and to get full credit the metric must match the human evaluation statistical significance (e.g. $p_{ij}^{m}\approx p_{ij}^{m}\approx 0.5$)
2.2 Addressing Metric Ties in PA
The fact that PA considers only binary wins/losses (e.g. the rounding in Equation 3) results in an interesting shortcoming in PA. There are (${{N\choose 2}}$ pairs of $N$ systems, and thus (${{N\choose 2}}+1$ distinct values that PA can take on ($0/{N\choose 2}$, $1/{N\choose 2}$, …, (${N\choose 2}/{N\choose 2}$). By way of example, WMT-22 en-de, with $N=14$ systems, (${{N\choose 2}}+1=92$.
However, metrics tend to perform better than a random baseline, so only the upper half of the range is actually useful (e.g. this leaves 46 distinct values for $N=14$ systems). In practice, most metrics perform substantially better than random, and we tend to compare many (e.g. >10) metrics. In practice, we find that this results in PA reporting the same score for several sets of metrics (see § 4.3).
By removing this rounding/quantization, SPA has no such issues.
3 Experimental Setup
3.1 Data
We conduct experiments on the data from the WMT Metrics shared tasks, years 2022 and 2023 Freitag et al. (2022, 2023). In particular, we use the primary language pairs where MQM judgments were collected. We use the MT Metrics Eval V2 toolkit [^1] https://github.com/google-research/mt-metrics-eval to retrieve official shared task scores.
We make the fairly arbitrary decision to compare all metrics, including non-primary metrics but excluding QE metrics (i.e. reference-free metrics) which provide segment-level scores.
In order to compute the statistical significance of comparisons between metrics metrics, we make the simplifying assumption that all system-level metrics are the average of their segment-level metric. This is not true for some metrics, including BLEU Papineni et al. (2002) and chrF Popović (2015). While it would be possible to re-compute BLEU and chrF for each subset, we average for simplicity. Our results are self consistent (i.e. we use the segment-level average in all cases) but will not accurately reflect the true performance of these metrics. To the best of our knowledge, this approximation is also made in recent WMT shared tasks.
3.2 $p$-value Computation
Figure 2: Final metric ranking stability when ablating the number of MT systems (and thus the number of total MQM judgments), measured as change in Pearson correlation coefficient (Pearson $r$) from the ranking computed on all MT systems. We find SPA to be more stable than PA in all cases.
Given a pair of translations of the same test set, by two different MT systems, each scored at the segment level by the same metric, we would like to know the statistical significance of the difference in the means of the segment-level scores of each MT system.
We use a permutation test Fisher (1935) to estimate the probability that our observed difference occurred randomly under the null hypothesis that the two translations have the same quality. A permutation test works by generating random splits of the data (ignoring the labels, i.e. which MT system that produced each translation) and computing the difference in metric score mean for these two sets of data. The $p$-value is computed by calculating the fraction of the time that this difference is greater than or equal to the mean difference in scores for the two translations of the test set.
Permutation tests are appealing because they don’t require any assumptions about the underlying distribution of the data. This fits our use case well because we do not want to assume anything about the distribution of segment-level scores of a metric, which can obviously change depending on the particular metric. [^2] Metric and human annotation distributions are both highly variable Lo et al. (2023b); Knowles and Lo (2024). Permutation tests instead have the assumption of exchangeability Pitman (1937); Draper et al. (1993); Good (2002)—that is, under the null hypothesis (e.g. the MT systems are equally good) the joint distribution of the observations is invariant under permutations of the data labels. To help ensure exchangeability, we perform permutations such that each split has exactly one translation of each test set sentence, commonly referred to as a paired permutation test Good (2013). We estimate each $p$-value from 1000 permutations.
3.3 $p$-value Speed Optimization
A naive implementation of the paired permutation test is not computationally prohibitive when computing $p$-values for all systems/metrics a single time, but it becomes a problem when we want to compute these values many times in order to estimate statistical significance of metric comparisons.
Experimentally, we find the main speed bottleneck to be generating the random permutations, so when estimating statistical significance of metric comparisons we cache a batch of permutations and use it for each pair of systems, on a per test-set basis. Additionally, by sharing permutations across system pairs, this allows us to pre-compute the contribution of each system to means of the random permutations, allowing computations to be linear instead of square in the number of MT systems. See our code [^3] https://github.com/thompsonb/mt-metrics-eval/blob/main/mt_metrics_eval/pairwise_paired_permutation_test.py for full implementation details. This gave us an >1000x speedup over the default implementation in Scipy Virtanen et al. (2020). [^4] Using the ‘permutation_test’ function from scipy, ‘permutation_type’ set to ‘samples’ and the ‘n_resamples’ set to 1000, each $p$-value takes around 40 milliseconds to compute on a laptop CPU. Our approach has no impact on a $p$-value for a single system-level comparison, but it does mean that the $p$-value for one pair of systems is no longer independent from the $p$-value for any other pair of systems. Given our use case (i.e., we are using these $p$-values as an approximate level of confidence for the system-level comparisons in the meta-metric formulation, as opposed to making any claims about the actual statistical significance of the system-level comparisons) we believe this lack of independence should be inconsequential.
4 Analysis
Meta-metric evaluation is challenging because there is no ground truth (i.e., we don’t know the true ranking of the metrics). Instead, we must make some assumptions about how an ideal metric will act. First, we study how sensitive the meta-metric results are when varying the number of MT systems and number of segments per MT system, with the assumption that lower sensitivity to the exact systems / segments used indicates a better meta-metric. Second, we examine whether PA indeed has the problem of ties that we hypothesized, and whether SPA fixes this issue. Finally, we test our hypothesis that the quantization/rounding in PA is effectively acting as additive random noise, and that SPA is effectively the same underlying meta-metric with the noise term removed. If this is true (and the magnitude of the noise does not dominate the underlying signal), we would expect SPA to produce a similar metric ranking to PA, but with increased statistical significance.
4.1 Ablation: Number of Systems
Each year, WMT and the associated metrics task collect and score many online and submitted MT systems. For an ideal meta-metric, the exact choice of MT systems would have little to no impact on the metric rankings. We perform an ablation on the number of MT systems being scored, keeping the number of annotations per system fixed. We then compute the correlation (as measured by Pearson’s $r$) between the meta-metric’s ranking of the ablations compared to that same meta-metric’s full ranking. This allows us to evaluate how sensitive the metric is to the exact selection of MT systems.
When ablating the number of MT systems (and keeping the number of annotations per system fixed), we find (see Figure 2) that SPA is more stable than PA across all MQM language pairs in the last two years of WMT metrics shared tasks.
4.2 Ablation: Sample Size
Since SPA relies on the pairwise $p$-values between MT systems, it is also natural to ask how SPA behaves when the number of available segments used for evaluating systems is small since it is harder to find statistical differences between systems with a smaller sample size. To answer this question, we calculated 95% confidence intervals for both PA and SPA values of two highly performant metrics—in particular, we considered xCOMET (Guerreiro et al., 2023) and MetricX-23 (Juraska et al., 2023)—on WMT23 using bootstrapping for various numbers of segments, thereby simulating scenarios with less human annotations but a fixed number of MT systems. The results are presented in § 4.2.
Figure 3: The 95% confidence intervals for SPA (blue) and PA (red) on Metric-X (top) and XCOMET (bottom) when varying the number of annotations per system. We find that SPA has a tighter confidence interval, and that the confidence interval shrinks to its full value with smaller sample sizes than PA.
When ablating the number of segments per MT system (and keeping the number of MT systems fixed), we find (see Figure 3) that SPA has tighter 95% confidence intervals than PA (shown on Metric-X and xCOMET), and that the confidence interval converges to its final value with smaller sample sizes than PA.
4.3 Ties
As discussed in § 2.2, the quantization noise in PA limits the number of distinct values it can assign to metrics. PA can only take on (${{N\choose 2}}+1$ distinct values. In practice, we find it tends to take on far fewer values. For example, in WMT 22 en-de PA could theoretically take on 92 distinct values, but because the metrics fall in a fairly narrow range of performance (between 62.6% for the worst metric and 81.3% for the best), the 21 metrics have only 11 distinct PA scores, with one 5-way PA tie and several 2- and 3-way PA ties (see Figure 4). Since SPA does not quantize each system comparison, it is able to assign any value to each metric, and is therefore potentially better able to distinguish between metrics.
Results for all language pairs are in Table 1. We find that on average, PA produces about half as many distinct values as there are metrics while SPA produces one unique value per unique metric.
4.4 Statistical Significance of Metric Comparisons
Just because a meta-metric produces two distinct scores for two different metrics does not mean that those metrics are statistically significantly different from each other. Ideally, a meta-metric would be able to produce statistically significant comparisons between any two distinct metrics. To explore the ability of PA and SPA to distinguish between metrics in a statistically significant way, we follow PERM-INPUTS Deutsch et al. (2021) to measure the statistical significance of the comparison between each metric. We follow recent shared tasks in greedily computing significance clusters, by starting with the highest scoring metric and assigning rank 1 to all metrics until we encounter the first metric that is statistically different from any previous metric so far. That metric is assigned rank 2, and the process repeats until all metrics have been assigned a rank. We echo the shared task organizers warning that this method can place two metrics that are statistically indistinguishable in different significance clusters (and does so multiple times, in the case of PA). Results are in § 4.4.
[…]
On average, SPA increases the number of statistically significant pairwise comparisons by 31% and the number of significance clusters by 40% compared to PA, while producing similar scores for each metric (see Table 1 for results summary, and Figure 4 for a visualization for WMT22 En-De results). This is consistent with our hypothesis that PA is effectively SPA with additive quantization noise.
Distinct Metric Significant Significant
Testset Language # MT # MT Values ($\uparrow$) Comparisons ($\uparrow$) Clusters ($\uparrow$)
Pairs Systems Metrics PA SPA Max PA SPA Max PA SPA Max
[…]
Table 1: Number of distinct values produced, number of statistically significant pairwise comparisons (p-value $<=0.05$), and number of statistical significance clusters for PA and SPA. We provide the best possible value for each category (Max) for comparison, but note that even an ideal meta-metric would likely not achieve this value due to some metrics being highly correlated with each other (e.g. due to training on the same data). : InstructScore and SEScoreX scores as returned by MT Metrics Eval V2 for WMT23 En-De are identical, causing an exact tie in both PA and SPA. We believe this is an error in MT Metrics Eval V2 but for posterity keep them as-is.
5 Historical Context and Related Work
WMT has run a machine translation evaluation since 2006 Koehn and Monz (2006). Since 2007 Callison-Burch et al. (2007), there has also been meta-evaluation of automatic metrics on the submitted translation systems. Here we summarize the rich 17 year history of system-level meta-evaluation at the WMT metrics shared tasks [^5] The WMT shared tasks have typically evaluated at both the system- and segment-level, but we focus on system-level meta-evaluation as it is most relevant to our work. and work related to and directly impacting the shared tasks, in order to demonstrate how our work fits into the historical context.
In the WMT 2007-2013 metrics evaluations Callison-Burch et al. (2007, 2008, 2009, 2010, 2011, 2012); Macháček and Bojar (2013) Spearman’s rank correlation coefficient $\rho$ was used for meta-evaluation of metrics. This was motivated by the fact that Spearman’s makes fewer assumptions about the data.
The WMT 2013 translation shared task Bojar et al. (2013) introduced system clusters (groups of systems that cannot be distinguished given the human judgments), and the 2013 metrics task Macháček and Bojar (2013) introduced empirical confidence of Spearman’s $\rho$ using bootstrap resampling. Since they were not able to resample on the submitted metrics, they only re-sampled human judgments. This iteration also discussed the fact that Spearman’s $\rho$ does not give partial credit. The penalty is equal for all wrong judgments, regardless of if the systems are close or far in quality. To compensate they present additional methods of analysis: Pearson’s, and correlation with systems’ clusters from the translation task Bojar et al. (2013). Those clusters were treated as ‘ranks with ties,’ and then correlation computed against Pearson’s and Pearson’s correlation against ‘fuzzy ranks’ (the average over ranks of all systems that are not significantly different in human quality).
In 2014, the metrics task Macháček and Bojar (2014) fully switched to Pearson’s $r$ from Spearman’s $\rho$. They also did bootstrap resampling to get empirical confidence intervals of system level correlations. This change to Pearson’s was due to the concerns pointed out in the previous year’s shared task, which had explored other meta-metrics.
The 2015 metrics task Stanojević et al. (2015) continued with Pearson’s $r$, and also presented analysis of Pearson’s $r$ vs Spearman’s $\rho$, and highlighted the instability of Spearman’s $\rho$ when MT systems are similar.
The 2016 metrics task Bojar et al. (2016) stuck with Pearson’s $r$, but changed the confidence to be the Williams test Williams (1959), as Graham and Baldwin (2014) had noted that this test is appropriate for dependent correlations.
The 2017 metrics task Bojar et al. (2017) kept Pearson’s $r$, and Williams test. They also added a pairwise significance test using Williams test. This continued in 2018 and 2019 Ma et al. (2018, 2019)
The 2020 metrics task Mathur et al. (2020b) continued to use Pearson’s, but also includes Kendall’s Tau for analysis. Kendall’s Tau is a closer match for the system ranking use case, since it is evaluating if the ordering of a pair of systems is the same as the human ordering. However it is does not take in to account the magnitude difference.
In 2021, the metrics task Freitag et al. (2021) switched to Pairwise accuracy Kocmi et al. (2021), motivated in part by the fact that MT system outliers had an outsized impact on Pearson correlation when it is used to rank MT Metrics Mathur et al. (2020a). The PERM-BOTH hypothesis test of Deutsch et al. (2021) was used to determine significance. 2021 and 2022 Freitag et al. (2022, 2023) follow.
In summary, the historical context of the WMT metric evaluations demonstrates that meta-evaluation is very challenging due to the numerous issues that must be simultaneously addressed, and underscores the pitfalls of making changes to meta-evaluation without considering the full set of ramifications. Most relevant to our work, it appears that the switch to pairwise accuracy in 2021 reduced the influence of outliers Mathur et al. (2020a) and (somewhat) aligned meta-evaluation with the standard use of comparing two systems with a metric, but it also reintroduced a problem that was first pointed out by Macháček and Bojar (2013) and more fully addressed by the change to Pearson’s $r$ from Spearman $\rho$ by Macháček and Bojar (2014): a disregard for the magnitude of differences. We address this issue by considering empirical confidence, which was first added by Macháček and Bojar (2013), and in the process we also better align meta-evaluation to the (more correct) usage of comparing two systems with a metric while also considering the statistical significance of the results.
5.1 Additional Connections to Prior Work
Graham and Liu (2016) proposed a method of sampling translations from every pair of competing MT systems, creating synthetic systems for scoring. Our work has clear similarities in that we create and score synthetic permutations, but differs in how those synthetic systems are used in the meta-metric formulation.
Mathur et al. (2020a) showed that MT system outliers had an outsized impact on Pearson correlation. In SPA, outliers impact is limited because $p$-values saturate at 0 or 1.
Knowles (2021) highlights that as WMT annotation protocols have shifted the original statistical assumptions, and questions the validity of the resulting protocols. Similarly, we show that shifts over the years have caused problems in meta-evaluation.
Lo et al. (2023a) investigated what magnitude of metric changes tend to be statistically significant. SPA uses statistical significance measures ($p$-values) directly, as opposed to the magnitude of metric differences (e.g. as in Pearson correlation).
Deutsch et al. (2023) demonstrated that principled tie handling is crucial when comparing MT metrics at the segment (e.g. sentence) level, because some metrics produced quantized scores that commonly create ties. SPA is system level (e.g. sentence level scores averaged over the entire test set), so exact ties are very unlikely. However, SPA can be seen as giving full credit for (statistical) ties, which is similar in spirit.
We show that Quantization is problematic in PA. Quantization in evaluation has proved problematic in other spaces as well—for example, Schaeffer et al. (2024) attributes the widely repeated claim that LLMs have emergent properties to quantization in evaluation.
6 Conclusions
We introduce a new meta-metric which we denote Soft Pairwise Accuracy, and show that it improves on Pairwise Accuracy in a number of ways, most notably that it is more stable when ablating the number of systems and annotations per system, and that it produces more statistically significant comparisons between metrics. We also discuss how SPA fits into and builds upon the nearly two decade history of meta-evaluation at WMT metric shared tasks.
Acknowledgments
This work would not have been possible without the WMT shared tasks and metrics task releasing their data. We are appreciative of helpful discussions with the WMT metrics task organizers and Rebecca Knowles.
Limitations
When computing $p$-values, we assume that system-level metric scores are the average of segment-level metrics scores. There is a line of recent work that seeks to incorporate contextual information into automatic metrics. Many such works still produce scores at the segment level (e.g. Vernikos et al., 2022; Hu et al., 2023; Agrawal et al., 2024) but others produce one score per window of a few sentences Raunak et al. (2024) or one score per paragraph Hu et al. (2023). Our method should still be applicable in such cases, but would require permuting windows or paragraphs instead of segments. Additionally, as previously noted, some metrics—notably BLEU Papineni et al. (2002) and chrF Popović (2015)—compute statistics at the segment level and combine them to create document-level scores. Again, permutations would still work but would require some modification. To the best of our knowledge, this issue is not limited to our work - it the same assumption is made in prior work computing statistical significance of metrics, including the WMT shared tasks Freitag et al. (2021, 2022, 2023) and Deutsch et al. (2021).
It is worth noting that the permutations in this work (as in prior works) are done on a single test set, and do not necessarily reflect variations in performance that could result from using the metrics in another domain. Prior work has shown that trained metrics are sensitive to a shift in domain relative to the data domain they were trained on Zouhar et al. (2024).